# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

### v3

@@ 4.1 Experimental Setup
CDAN [42] 54.48 53.87 42.88 36.21 46.86
MemSAC [25] 53.02 54.37 42.05 38.33 46.94
ToAlign [70] 55.67 55.92 42.32 38.40 48.08
MDD [77] 51.57 50.73 42.54 39.23 46.02
DALN [7] 55.36 55.77 41.06 40.41 48.15
PMTrans [79] 56.76 57.60 46.18 40.33 50.22
@@
TextMatch [27] 49.68 54.82 53.06 50.11 51.92
nGramMatch [27] 49.53 51.02 51.70 49.87 50.93
LaGTran [27] 63.67 64.16 56.14 57.02 60.24
@@
Improvement +3.72 +5.81 +5.01 +6.49 +5.26
Table 2: Results on GeoImnet [26] and GeoPlaces [26] with 4 adaptation scenarios. Best results are in bold, whereas second best results are underlined. All methods use the ViT-B [12] backbone.
4.2 Results
We benchmark our model against various language-free domain adaptation methods that have reported results on the considered datasets [79, 70, 25, 77, 7, 75]. Besides classic UDA methods, we also compare with previous approaches using language guidance [27, 56, 16, 31, 36]. Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
DomainNet. Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P). LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison. LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for the Swin-B backbone. For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$. Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-specific prompting, unlike previous models [31, 36, 16].
[…]
These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of LAGUNA with respect to both language-free and language-guided UDA approaches.
@@ 4.2 Results
TextMatch [27] 10.36 13.57 11.97
nGramMatch [27] 11.50 15.46 13.98
LaGTran [27] 12.34 30.76 21.55
@@
Table 3: Results on the Ego2Exo benchmark from EgoExo-4D [20]. Best results are in bold, whereas second best results are underlined. All methods use official pre-extracted Omnivore-L features [17, 20] except EgoVLP [37] and LaVILA [78], which use Timesformer-B [3] backbone.
@@ 4.3 Ablation Study
We analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify design choices. We investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces. More ablations are reported in the Supplementary Material.
@@
Methodological Elements. Table 4, ablates the main methodological elements of LAGUNA on GeoImnet. We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ Eq. (6)), domain-agnostic reference anchors $\mathcal{A}$, learnable anchors $\mathcal{A}_{t/s}$, cross-domain attention (CD Attn. - Eq. (9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq. (12)). We then progressively add these elements (settings (1)-(5)) until we reach LAGUNA. For all settings, we report the relative improvement (w.r.t. previous row) and the absolute one (w.r.t. first row).
@@
Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27]. In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity. This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
@@
Relative alignment In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq. 6). The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one.
@@
Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1).
@@
Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to a of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1).
Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to (1). Together, these improvements show the effectiveness of LAGUNA, highlighting the benefits of learning domain-specific visual representations aligned to a reference language structure. Language models for reference structures. Fig. 3 reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer [57], CLIP [56], and BERT [61]), together with the average accuracy obtained keeping stages 2-3 fixed. Notably, accuracy follows the ability of models to distinguish distinct classes, with low off-diagonal values indicating more discriminative models. While we achieve best results with SentenceTransformer, LAGUNA’s performance is robust to different reference structures, maintaining SOTA results also when the same language model is used across all stages (i.e., BERT).
Figure 3: Similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies.
Ratio of target samples. LAGUNA demonstrates strong generalization capabilities even with limited target domain data (Table 5). We progressively increase the number of randomly selected pseudo-labeled target data used for training (10%-75%) and observe the impact on performance. Notably, LAGUNA achieves high accuracy even with only $10\%$ of the samples and attains SOTA results with just $20\%$ (compared with  [27]).
Dataset 10% 20% 30% 50% 75% 100% [27]
GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68 63.92
GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33 56.58
Avg. 58.38 60.64 61.58 63.22 63.72 65.40 60.24
Table 5: Average accuracy with different percentages of target data compared with previous SOTA LaGTran [27]. LAGUNA to obtains SOTA accuracy with $20\%$ (in green).
Alignment in Absolute and Realtive Coordinates. Figure 4 shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where source is USA and target is Asia. Notably, features are separated in absolute coordinates, while they blend together (low MMD value) in relative coordinates thanks to the structure-preserving loss $\mathcal{L}_{S}$ and domain-specific learnable anchors. This highlights that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
@@
Figure 4: t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD [28] scores. Relative coordinates (right) are better aligned than absolute (left).
Figure 5: Ablation with BLIP-1 and BLIP-2 generated captions.
@@
Computational Overhead. While LAGUNA involves three stages and different components, its training is quite efficient: Stage 1 needs no training, and embedding 600 class labels takes 1-2 minutes ($\sim$4GB GPU memory). On GeoImNet, with a V100, Stage 2 requires only 9 minutes per epoch, while Stage 3 adds $<$2 minutes per epoch, which is minor considering the baseline’s 65 minutes per epoch. At test time, LAGUNA only requires the visual encoder, the learned anchors, the light cross-attention layer and the final MLP classifier, bringing no noticeable overhead.
@@ 5 Conclusions
We introduced LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation. LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics. By conditioning the classifier to adhere to this structure, LAGUNA encourages structural similarity across domain-specific latent spaces, retaining unique features for improved classification. This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment. Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods.
Supplementary Material
@@ Appendix A Introduction
In this document, we report additional ablation studies regarding the choice of language models and the structural guidance of the target domain. Moreover, we provide a more detailed explanation of the motivation behind our regularization loss $\mathcal{L}_{Reg}$ in Eq. (8). Finally, we attach a Python file, LAGUNA.py, demonstrating the implementation of LAGUNA in PyTorch. We will release the full code and configurations upon acceptance.
Appendix B Ablation on Language Models
In addition to the language model used for defining the reference structure through $\mathcal{A}$ in Stage 1, LAGUNA employs language models in Stage 2 for encoding image/video captions trained for text classification and capturing semantic structure. This model is then used to generate pseudo-labels and serve as structure guidance for the target domain in Stage 3. We ablate the choice of language model in these Stages, comparing LAGUNA’s performance when employing SentenceTransformer [57], CLIP [56], or BERT [61] on GeoImnet dataset in two settings: (1) using SentenceTransformer-defined $\mathcal{A}$ (Table 6) as it can better model semantic relationships (recall Fig. 3 from the main manuscript) and (2) defining $\mathcal{A}$ with the same language model as the one chosen for training (Table 7).
@@ Appendix B Ablation on Language Models
Appendix C Ablation on structure guidance for target domain in Stage-3
@@
Appendix D Regularization Loss Extended

### v4

@@ 4.2 Results
CDAN (Long et al., 2018) 54.48 53.87 42.88 36.21 46.86
MemSAC (Kalluri et al., 2022) 53.02 54.37 42.05 38.33 46.94
ToAlign (Wei et al., 2021) 55.67 55.92 42.32 38.40 48.08
MDD (Zhang et al., 2019) 51.57 50.73 42.54 39.23 46.02
DALN (Chen et al., 2022) 55.36 55.77 41.06 40.41 48.15
PMTrans (Zhu et al., 2023) 56.76 57.60 46.18 40.33 50.22
@@
TextMatch (Kalluri et al., 2024) 49.68 54.82 53.06 50.11 51.92
nGramMatch (Kalluri et al., 2024) 49.53 51.02 51.70 49.87 50.93
LaGTran (Kalluri et al., 2024) 63.67 64.16 56.14 57.02 60.24
@@
Improvement +3.72 +5.81 +5.01 +6.49 +5.16
Table 2: Results on GeoImnet and GeoPlaces with 4 adaptation scenarios using ViT-B. Best results are in bold, whereas second-best are underlined.
@@
TextMatch (Kalluri et al., 2024) 10.36 13.57 11.97
nGramMatch (Kalluri et al., 2024) 11.50 15.46 13.98
LaGTran (Kalluri et al., 2024) 12.34 30.76 21.55
@@
Table 3: Results on the Ego2Exo benchmark. Best results are in bold, second best are underlined. All methods use Omnivore-L features except EgoVLP and LaVILA.
@@ 4.3 Ablation Study
In this section, we conduct comprehensive ablation studies across multiple dimensions of LAGUNA: 1) methodological component analysis, 2) language model assessment for reference structure generation, 3) qualitative comparison of relative vs. absolute alignment in LAGUNA’s features, 4) comparison with zero-shot MLLMs, 5) caption quality impact analysis, and 6) universal domain adaptation performance evaluation. Additional complementary ablation studies are provided in the supplementary material.
@@
Methodological Elements. Table 4, ablates the main methodological elements of LAGUNA on GeoImnet. We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ - Eq. (5)), domain-agnostic reference anchors $\mathcal{A}$, learnable anchors $\mathcal{A}_{t/s}$, cross-domain attention (CD Attn. - Eq. (8), (9)), and the regularization loss ($\mathcal{L}_{Reg}$ - Eq. (7)). We then progressively add these elements (settings (1)-(5)) until we reach LAGUNA. For all settings, we report the relative improvement (w.r.t. previous row) and the absolute one (w.r.t. first row).
@@
Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran. In (2), we add $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity. Setting (2) improves (1) by +0.87+0.87, suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
@@
Relative alignment In (3), we extend (2) with domain-independent learnable anchors ($\checkmark*$ in Table 4, indicating that weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the $\mathcal{L}_{S}$ loss (Eq. 5). The introduction of learnable anchors enables the relative alignment of the visual domain with $\mathcal{A}$, leading to gains of +1.32+1.32 w.r.t (2) and +2.10+2.10 over (1), highlighting the benefits of the proposed relative alignment.
@@
Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of +1.77+1.77 w.r.t (3) and a robust +3.87+3.87 w.r.t the baseline (1).
@@
Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to+0.44+0.44 w.r.t (4) and an overall improvement of +4.31+4.31 w.r.t (1).
Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of +1.16+1.16 w.r.t (5) and an overall gain of +5.47+5.47 w.r.t (1). These improvements show the efficacy of LAGUNA and the benefits of aligning visual representations to a reference structure.
@@
Figure 3: In a), similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies. In b), t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD scores in Relative (right) and absolute (left) spaces.
Language models for reference structures. Fig. 3 a) reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer, CLIP, and BERT), together with the average accuracy obtained keeping stages 2-3 fixed. Notably, accuracy follows the ability of models to distinguish classes, with low off-diagonal values indicating more discriminative models. While we achieve best results with SentenceTransformer, LAGUNA is robust to different reference structures, maintaining SOTA results.
Absolute vs Relative Alignment. Figure 3 b) shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where the source is USA. Notably, features are separated in absolute coordinates, while they blend together (low MMD) in relative ones thanks to $\mathcal{L}_{S}$ and domain-specific learnable anchors. This highlights that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
Comparison with ’Zero-Shot’ MLLMs. Despite recent advances in MLLMs with billions of parameters trained on web-scale data, our experiments reveal that LAGUNA significantly outperforms them in challenging domain adaptation scenarios. Testing on two benchmarks requiring fine-grained spatial and temporal discrimination (GeoNet and Ego2Exo, Tab 5&6), we compare against five SOTA 7-8B MLLMs. LAGUNA achieved +11.84% higher accuracy on GeoNet and +2.78% in mean per-class accuracy on Ego2Exo. More importantly, in the Ego2Exo benchmark, we surpass InternVL-8B, which is trained in egocentric data and holds the state-of-the-art in many egocentric benchmarks. As a final remark, LAGUNA is approximately 100× smaller than these MLLMs, demonstrating that tailored models are still better at challenging UDA tasks.
Model Geolmnet GeoPlaces Avg.
@@
LLaVA-Next 7B Li et al. (2024b) 22.88 24.21 40.41 42.33 32.46
LLaVA-OneVision 7B Li et al. (2024a) 27.07 29.87 46.29 49.23 38.12
Qwen2.5-VL 7B Bai et al. (2025) 52.98 56.34 50.85 54.11 53.56
InternVL3 8B Zhu et al. (2025) 38.95 41.58 53.66 52.91 46.75
PLM Cho et al. (2025) 19.93 23.96 35.58 42.61 30.52
LAGUNA 67.39 69.97 61.15 63.51 65.40
Improvement +14.41 +13.63 +7.49 +9.40 +11.84
Table 5: Comparison on GeoImnet and GeoPlaces against zero-shot MLLMs.
[…]
LLaVA-Next 7B Li et al. (2024b) 4.28 4.95 4.62
LLaVA-OneVision 7B Li et al. (2024a) 5.57 10.18 7.88
Qwen2.5-VL 7B Bai et al. (2025) 11.46 20.39 15.93
InternVL3 8B Zhu et al. (2025) 13.07 28.34 20.71
PLM Cho et al. (2025) 22.54 37.16 29.85
LAGUNA 13.52 33.45 23.49
Improvement +0.45 +5.11 +2.78
Table 6: Comparison on Ego2Exo against zero-shot MLLMs. PLMs training set includes EgoExo-4D.
[…]
LLaVA-OneVision 7B Li et al. (2024a) 82.20 74.10 78.15
Qwen2.5-VL 7B Bai et al. (2025) 82.35 74.26 78.31
LAGUNA (LLaVA captions) 82.51 76.20 79.35
LAGUNA (Qwen2.5 captions) 82.77 76.19 79.48
Table 7: Ablation on the quality of generated captions with different MLLMs on sketch and clipart domains from DomaiNet.
Model H-Score
Baseline Kalluri et al. (2024) 50.20
UniDA You et al. (2019) 33.39
DANCE Saito et al. (2020) 51.75
OVANet Saito and Saenko (2021) 47.26
LaGTran Kalluri et al. (2024) 61.19
LAGUNA 64.41
Table 8: Experiment on GeoUniDA for universal domain adaptation.
Caption Quality. Table 1 presents LAGUNA results using BLIP-2-generated captions for consistency and fair comparison with previous works. In Table 7, we investigate how caption quality affects LAGUNA’s performance by generating captions from two SOTA MLLMs: LLaVA-OneVision 7B and Qwen2.5-VL 7B, and evaluating LAGUNA on DomainNet’s sketch and clipart domains, comparing its performance against the same MLLMs. Results show that LAGUNA benefits from higher-quality captioning models and offers an alternative that is 100x smaller while achieving better performance on specific domains. The supplementary material provides additional experiments demonstrating: 1) LAGUNA’s robustness to low-quality captions and 2) its requirement for only a few target captions to achieve SOTA performance, reducing the one-time computational cost of automatic annotation with MLLMs. These findings make LAGUNA well-suited for real-world deployment in constrained environments where billion-parameter models are impractical.
Universal Domain Adaptation. Finally, we assess LAGUNA’s performance in open-world domain transfer scenarios using the GeoUniDA dataset Kalluri et al. (2023), which, in contrast with traditional domain adaptation benchmarks, evaluates universal domain adaptation across domains, including both unique and shared semantic classes. Following OVANet Saito and Saenko (2021), we employ the H-score metric that balances closed-set and open-set accuracies through their harmonic mean. Notably, LAGUNA exhibits strong performance in this challenging setting, validating the effectiveness of our method.
@@ 5 Conclusions
We introduce LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation. LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics. By conditioning the classifier to adhere to this structure, LAGUNA encourages structural similarity across domain-specific latent spaces, retaining unique features for improved classification. This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment. Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods, highlighting the importance of structural alignment and language-guided learning in domain adaptation. Future work could explore extending LAGUNA to multi-modal adaptation scenarios.
Appendix LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces
Figure 4: LAGUNA’s accuracy with different quantities of pseudo-labeled data (target samples with captions).
@@ Appendix A Introduction
This supplementary material presents additional ablation studies complementing those in the main manuscript. We first examine the number of captions LAGUNA requires to achieve SOTA performance. Second, we extend caption quality experiments by testing LAGUNA’s robustness with lower-quality captions. Third, we ablate language model choices and target domain structural guidance. Fourth, we provide detailed motivation for our regularization loss $\mathcal{L}_{Reg}$ in Eq. (8). Full code and implementation will be released upon acceptance. In addition to LAGUNA’s code, we will also provide scripts to reproduce all the MLLMs results reported in this work.
Appendix B Ablation on the ratio of target samples.
Our approach relies on image captions, whether generated or sourced elsewhere. A potential concern with LAGUNA is that automatic annotation of large datasets can be tricky to obtain in full or even time-consuming. While the costs of obtaining these captions are minimal compared to manual annotation, we investigate the quantity of captions LAGUNA requires for SOTA performance in Fig. 4. We progressively increase randomly selected pseudo-labeled training data from 10% to 75% and observe the performance impact. Notably, LAGUNA achieves high accuracy with only 10% of samples and reaches SOTA results with just 20%. This efficiency highlights the benefits of our structure-driven approach, which organizes sample projections within domain-specific spaces according to a reference geometric structure.
Appendix C Ablation on captioning quality.
[…]
Model S $\rightarrow$ C C $\rightarrow$ S
BLIP-1 Li et al. (2022) 69.9 60.1
BLIP-2 Li et al. (2023) 72.0 64.1
LAGUNA (BLIP-1) 78.1 71.3
LAGUNA (BLIP-2) 80.2 73.5
Table 9: Ablation on caption quality with BLIP-1&2 captioning models. We consider sketch and clipart domains from DomainNet. The best result is in bold, and the second best is underlined.
Appendix D Ablation on Language Models
In addition to the language model used for defining the reference structure through $\mathcal{A}$ in Stage 1, LAGUNA employs language models in Stage 2 for encoding image/video captions trained for text classification and capturing semantic structure. This model is then used to generate pseudo-labels and serve as structure guidance for the target domain in Stage 3. We ablate the choice of language model in these Stages, comparing LAGUNA’s performance when employing SentenceTransformer Reimers and Gurevych (2019), CLIP Radford et al. (2021), or BERT Sanh et al. (2019) on GeoImnet dataset in two settings: (1) using SentenceTransformer-defined $\mathcal{A}$ (Table 10) as it can better model semantic relationships (recall Fig. 3 from the main manuscript) and (2) defining $\mathcal{A}$ with the same language model as the one chosen for training (Table 11).
@@ Appendix D Ablation on Language Models
Appendix E Ablation on structure guidance for target domain in Stage-3
@@
Appendix F Regularization Loss Extended
@@
Appendix G Use of generative AI
According to the ICLR 2026 policy regarding the transparency of generative AI use in submitted works, we declare that generative AI tools were employed solely for language enhancement and correction purposes in this manuscript. All content, including technical contributions, experimental design, analysis, and conclusions, represents original work by the authors. We take full responsibility for every word reported in this paper, as each statement has been carefully reviewed, verified, and deliberately selected to accurately represent our research findings and contributions. The use of AI assistance was limited exclusively to improving grammatical clarity and linguistic precision, without influencing the scientific content or integrity of our work.